12) 5*3-41a dx 15) Rewrite 14) szlo0 Vandx . dx with the correct limits. DO NOT EVALUTATE,
Answer #1

Take u=x-19 limit x=2 x = 19 duada u=-17 = 0 du -la - 60-80) . 12-19 Umit Take u=2-19 duada 12=19 uso 2=100 u=81 a+* a + b =D-da. - 2 al it 2=2 ; 2 it & an undebised point with in boundaries. exist b, acbce, f (b) = undebind T b=2 n= (a) dnt + f(n) d